OpenStudy (anonymous):

Use addition or substitution to find the value of x for this set of equations. 4x+13y=40 4x+3y=-40

terenzreignz (terenzreignz):

You can subtract the second expression form the first.

OpenStudy (anonymous):

i don't understand how would i do that?

OpenStudy (lgbasallote):

4x + 13y = 40 4x + 4y = -40 use subtraction to cancel the x 4x + 13y = 40 4x + 3y = -40 ========= 0 + 10y = 80 10y = 80 y = 80/10 y = 8 now use this to solve for x 4x + 3y = -40 4x + 3(8) = -40 4x + 24 = -40 4x = -40 - 24 4x = -64 x = -64/4 x = -16 do you get that @hispanicgal5

OpenStudy (anonymous):

ohhhh i get it thank you! @lgbasallote

OpenStudy (anonymous):

Use addition or substitution to find the value of x for this set of equations. 12x - 5y = 30 y = 2x - 6

OpenStudy (anonymous):

put your y in and solve for x 12x-5(2x-6)=30 12x-10x-30=30 2x=60 x=30 now put that in for x in the equation y=2x-6 and you have your answer